# State-Constrained Offline Reinforcement Learning

### v1

@@ Appendix C Implementation details
This means that the maximum state dimension model prediction error must be within $10\%$ of the true state range. Due to this we also normalise all states which gives more accurate model prediction and is done in the same way as TD3+BC [21], i.e let $s_{i}$ be the $i$th feature of state $s$
@@
For the MuJoCo locomotion tasks we evaluate our method over $5$ seeds each with $10$ evaluation trajectories; whereas for the Antmaze tasks we also evaluate over $5$ seeds but with $100$ evaluation trajectories. Just like TD3+BC [21], we scale our hyperparameter by the average Q-value across the minibatch
@@
The actor, critic and reward model are represented as neural networks with two hidden layers of size $256$ and ReLU activation. They are trained using the ADAM optimiser [30] and have learning rates $43e-$ 4, the actor also has a cosine scheduler. We use an ensemble of $4$ critic networks and take the minimum value across the networks. Also we use soft parameter updates for the target critic network with parameter $\tau=0.005$, and we use a discount factor of $\gamma=0.99$. For the locomotion tasks we use a shared target value to update the critic towards, whereas for the Antmaze tasks we use independent target values for each critic value. Both the inverse and forward dynamics models are represented as neural networks with three hidden layers of size $256$ and ReLU activation. They are trained using the ADAM optimiser with a learning rate of $34e-$ 3 and a batch size of $256$. We use an ensemble of $7$ forward models and $3$ inverse models and then take a final prediction as an average across the ensemble. Also our forward model predicts the state difference rather than the next state directly, which improves model prediction performance.
To attain the results for BCQ, we re-implemented the algorithm from the paper [22] and trained on the version 2 datasets. Following the same practises as StaCQ we evaluate BCQ over 5 seeds each with 10 evaluations on the MuJoCo locomotion tasks and $5$ seeds each with $100$ evaluations on the Antmaze tasks. All other results in Table 1 were obtained from the original authors’ papers. Our experiments were performed with a single GeForce GTX 3090 GPU and an Intel Core i9-11900K CPU at 3.50GHz. Each run takes on average $2.5$ hours, where models are pre-trained and state reachability is given (the same models and reachability lists are given to each run). We provide $5$ runs for each dataset ($18$) which gives a total run time of $225$ hours.
@@ Appendix D One-step method
StaCQ, Algorithm 1, that we have introduced in this paper is an actor-critic method that learns the QS-value and policy together. Alternatively, the QS-value can be learned directly from the data, in an on-policy fashion, then a policy can be extracted directly from this on-policy QS-value. These approaches are known as one-step methods [8]. In this section we adapt StaCQ into a one-step method.
D.1 Estimating QS-values
So that the QS-values are learned on-policy while still taking advantage of the state-constrained framework we introduce a small modification to Eq. (6). Since the pair $(s^{\prime},s^{\prime\prime})$ is unseen in the dataset, we use the following approximation:
@@ D.1 Estimating QS-values
to replace $\max_{\begin{subarray}{c}s^{\prime\prime}s.t.s^{\prime\prime}\in\mathcal{D}\\ s^{\prime\prime}\in\mathcal{SR}_{\mathcal{M}}(s^{\prime})\end{subarray}}Q(s^{\prime},s^{\prime\prime})$. This adjustment ensures that the QS-value is only evaluated on explicit state-next-state pairs, thereby avoiding OOD $(s,s^{\prime})$-pairs. Although this approach diverges slightly from the theoretical method, where QS-value updates are performed on every pair, it provides a practical solution.
Using Eq. (29), the approximation of state reachability and the reward model Eq. (12), the on-policy state-constrained QS-values can be refined by reducing the MSE between the target and the actual QS-values. The target is determined by identifying the maximum value across reachable states:
@@ D.2 Policy Extraction Step
Eq. (30) gives a one-step optimal state-constrained QS-value. However, this equation alone does not produce an optimal action. To determine the optimal action, we need to add a policy extraction step. We use the same policy extraction method as StaCQ, a state behaviour cloning regularised policy update similar to TD3+BC [21]:
@@
$\hat{s}^{\prime}=\arg\max_{\begin{subarray}{c}s^{\prime}\mathrm{s.t.}(s^{\prime},s^{\prime\prime})\in\mathcal{D}\\ s^{\prime}\in\widehat{\mathcal{SR}}_{\mathcal{M}}(s)\end{subarray}}\{r_{\omega_{3}}(s,s^{\prime})+\gamma Q_{\theta}(s^{\prime},s^{\prime\prime})\},$ (32)
@@
3:  Pre-train $f_{\omega_{1}}$ and $I_{\omega_{2}}$ using Eqs.(8) and (9)
@@
6:     Optimise reward function according to Eq. (12)
7:     Optimise QS-value according to Eq. (30)
@@ D.3 One step method implementation details
In the original OneStepRL paper [8], they evaluate their method over $10$ seeds where the $\lambda$ hyperparameter has been tuned over the first $3$ seeds and then evaluated with the $\lambda$ fixed for the remaining $7$. However as we want a consistent hyperparameter for each environment we choose $\alpha=\{0.1,1.0,5.0\}$ for Hopper, Walker2d and Halfcheetah respectively and for the Antmaze tasks we choose $\alpha=\{10,40,100\}$ for the umaze, medium and large environments respectively. The one step method also uses a single critic, we found that increasing the number of critics deteriorated results. All other implementation details are the same as StaCQ, Appendix C.

### v2

@@ Appendix C Implementation details
This means that the maximum state dimension model prediction error must be within $10\%$ of the true state range. Due to this we also normalise all states which gives more accurate model prediction and is done in the same way as TD3+BC (Fujimoto & Gu, 2021), i.e let $s_{i}$ be the $i$th feature of state $s$
@@
For the MuJoCo locomotion tasks we evaluate our method over $5$ seeds each with $10$ evaluation trajectories; whereas for the Antmaze tasks we also evaluate over $5$ seeds but with $100$ evaluation trajectories. Just like TD3+BC (Fujimoto & Gu, 2021), we scale our hyperparameter by the average Q-value across the minibatch
@@
The actor, critic and reward model are represented as neural networks with two hidden layers of size $256$ and ReLU activation. They are trained using the ADAM optimiser (Kingma & Ba, 2014) and have learning rates $43e-$ 4, the actor also has a cosine scheduler. We use an ensemble of $4$ critic networks and take the minimum value across the networks. Also we use soft parameter updates for the target critic network with parameter $\tau=0.005$, and we use a discount factor of $\gamma=0.99$. For the locomotion tasks we use a shared target value to update the critic towards, whereas for the Antmaze tasks we use independent target values for each critic value. Both the inverse and forward dynamics models are represented as neural networks with three hidden layers of size $256$ and ReLU activation. They are trained using the ADAM optimiser with a learning rate of $34e-$ 3 and a batch size of $256$. We use an ensemble of $7$ forward models and $3$ inverse models and then take a final prediction as an average across the ensemble. Also our forward model predicts the state difference rather than the next state directly, which improves model prediction performance.
To attain the results for BCQ, we re-implemented the algorithm from the paper (Fujimoto et al., 2019) and trained on the version 2 datasets. Following the same practises as StaCQ we evaluate BCQ over 5 seeds each with 10 evaluations on the MuJoCo locomotion tasks and $5$ seeds each with $100$ evaluations on the Antmaze tasks. All other results in Table 1 were obtained from the original authors’ papers. Our experiments were performed with a single GeForce GTX 3090 GPU and an Intel Core i9-11900K CPU at 3.50GHz. Each run takes on average $2.5$ hours, where models are pre-trained and state reachability is given (the same models and reachability lists are given to each run). We provide $5$ runs for each dataset ($18$) which gives a total run time of $225$ hours.
@@ Appendix D One-step method
StaCQ, Algorithm 1, that we have introduced in this paper is an actor-critic method that learns the QSS-value and policy together. Alternatively, the QSS-value can be learned directly from the data, in an on-policy fashion, then a policy can be extracted directly from this on-policy QSS-value. These approaches are known as one-step methods (Brandfonbrener et al., 2021). In this section we adapt StaCQ into a one-step method.
D.1 Estimating QSS-values
So that the QSS-values are learned on-policy while still taking advantage of the state-constrained framework we introduce a small modification to Eq. equation 6. Since the pair $(s^{\prime},s^{\prime\prime})$ is unseen in the dataset, we use the following approximation:
@@ D.1 Estimating QSS-values
to replace $\max_{\begin{subarray}{c}s^{\prime\prime}s.t.s^{\prime\prime}\in\mathcal{D}\\ s^{\prime\prime}\in\mathcal{SR}_{\mathcal{M}}(s^{\prime})\end{subarray}}Q(s^{\prime},s^{\prime\prime})$. This adjustment ensures that the QSS-value is only evaluated on explicit state-next-state pairs, thereby avoiding OOD $(s,s^{\prime})$-pairs. Although this approach diverges slightly from the theoretical method, where QSS-value updates are performed on every pair, it provides a practical solution.
Using Eq. equation 30, the approximation of state reachability and the reward model Eq. equation 13, the on-policy state-constrained QSS-values can be refined by reducing the MSE between the target and the actual QSS-values. The target is determined by identifying the maximum value across reachable states:
@@ D.2 Policy extraction step
Eq. equation 31 gives a one-step optimal state-constrained QSS-value. However, this equation alone does not produce an optimal action. To determine the optimal action, we need to add a policy extraction step. We use the same policy extraction method as StaCQ, a state behaviour cloning regularised policy update similar to TD3+BC (Fujimoto & Gu, 2021):
@@
$\hat{s}^{\prime}=\operatorname*{arg\,max}_{\begin{subarray}{c}s^{\prime}\mathrm{s.t.}(s^{\prime},s^{\prime\prime})\in\mathcal{D}\\ s^{\prime}\in\widehat{\mathcal{SR}}_{\mathcal{M}}(s)\end{subarray}}\{r_{\omega_{3}}(s,s^{\prime})+\gamma Q_{\theta}(s^{\prime},s^{\prime\prime})\},$ (33)
@@
3:  Pre-train $f_{\omega_{1}}$ and $I_{\omega_{2}}$ using Eqs.equation 9 and equation 10
@@
6:     Optimise reward function according to Eq. equation 13
7:     Optimise QSS-value according to Eq. equation 31
@@ D.3 One step method implementation details
In the original OneStepRL paper (Brandfonbrener et al., 2021), they evaluate their method over $10$ seeds where the $\lambda$ hyperparameter has been tuned over the first $3$ seeds and then evaluated with the $\lambda$ fixed for the remaining $7$. However as we want a consistent hyperparameter for each environment we choose $\alpha=\{0.1,1.0,5.0\}$ for Hopper, Walker2d and Halfcheetah respectively and for the Antmaze tasks we choose $\alpha=\{10,40,100\}$ for the umaze, medium and large environments respectively. The one step method also uses a single critic, we found that increasing the number of critics deteriorated results. All other implementation details are the same as StaCQ, Appendix C.
Appendix E Consistent hyperparameter experiment
The single StaCQ hyperparameter, $\lambda$ in the policy update, is tuned per environment (for example, using the same $\lambda$ for all Hopper datasets). This ensures StaCQ is not over-tuned to each dataset while still producing realistic, high-quality results. Most state-of-the-art methods have different hyperparameters for each dataset (Yu et al., 2020; Kidambi et al., 2020; Brandfonbrener et al., 2021; Wang et al., 2021; Ghasemipour et al., 2022; Rigter et al., 2022; Xu et al., 2022); this is not the case for StaCQ. However, TD3+BC and IQL both use a consistent hyperparameter across all datasets. TD3+BC uses an ensemble of two critics, whereas StaCQ has an ensemble of four. To compare fairly with these methods, we perform two comparisons. First, we fix StaCQ’s hyperparameter across all locomotion tasks. Second, we tune TD3+BC’s hyperparameter for each environment and supplement it with an ensemble of four critics. Table 2 shows these comparisons.
Fixed hyperparameter Per environment tuning & 4 critics
IQL TD3+BC StaCQ TD3+BC StaCQ
Random Halfcheetah - $11.0$ $\mathbf{15.8}$ $18.0\pm 1.1$ $\mathbf{24.3\pm 1.3}$
Hopper - $\mathbf{8.5}$ $\mathbf{8.6}$ $7.6\pm 0.7$ $\mathbf{17.5\pm 12.8}$
Walker2d - $1.8$ $\mathbf{3.5}$ $0.9\pm 0.2$ $\mathbf{4.4\pm 6.4}$
Med Halfcheetah $47.4$ $48.3$ $\mathbf{50.7}$ $49.0\pm 0.3$ $\mathbf{57.6\pm 0.6}$
Hopper $66.3$ $59.3$ $\mathbf{100.1}$ $61.5\pm 7.7$ $\mathbf{100.2\pm 3.0}$
Walker2d $78.3$ $83.7$ $\mathbf{88.6}$ $83.6\pm 4.8$ $\mathbf{92.2\pm 3.0}$
Med-Rep Halfcheetah $44.2$ $\mathbf{44.6}$ $\mathbf{46.5}$ $44.6\pm 0.7$ $\mathbf{52.2\pm 0.5}$
Hopper $94.7$ $60.9$ $\mathbf{98.7}$ $61.5\pm 23.8$ $\mathbf{99.1\pm 1.3}$
Walker2d $73.9$ $\mathbf{81.8}$ $\mathbf{85.5}$ $\mathbf{84.0\pm 3.0}$ $\mathbf{87.2\pm 7.3}$
Med-Exp Halfcheetah $86.7$ $\mathbf{90.7}$ $\mathbf{92.4}$ $90.9\pm 2.2$ $\mathbf{96.4\pm 2.9}$
Hopper $91.5$ $\mathbf{98.0}$ $\mathbf{94.0}$ $104.5\pm 11.8$ $\mathbf{111.9\pm 0.2}$
Walker2d $\mathbf{109.6}$ $\mathbf{110.1}$ $\mathbf{114.2}$ $\mathbf{110.6\pm 0.5}$ $\mathbf{116.2\pm 1.5}$
Total $692.6$ $698.7$ $\mathbf{798.6}$ $716.7$ $\mathbf{859.2}$
Table 2: Average normalised scores for StaCQ, TD3+BC and IQL with consistent hyperparameter tuning strategies. The left hand side compares StaCQ with IQL and TD3+BC where there is a consistent hyperparameter across all datasets. The right hand side compares StaCQ with TD3+BC with $4$ critics and hyperparameters tuned per dataset. The bold scores are within the $95\%$ of the highest performing method in each dataset. Scores represent mean over $5$ seeds of $10$ evaluation $\pm$ the standard deviation.
From Table 2, we can see that with consistent hyperparameter tuning strategies StaCQ still vastly outperforms TD3+BC (and IQL). This now provides a clear comparison showing how more Q-value evaluations through the state reachability metric allows for higher-quality policies than the batch-constrained counterparts.
Appendix F Ablation study: exploring the effect of the reachability metric
In this section, we explore the effect of changing the reachability metric for StaCQ. That is, for fixed datasets, Walker2d-Medium and Walker2d-MediumExpert, we explore how the change of reachability norm and threshold for reachable states, $\epsilon$, effects the average normalised score.
Table 3 shows the comparison of the reachability metric with different thresholds. For the $L^{1}$-norm with $\epsilon=0.01\text{ and }0.05$ and $l^{\infty}$-norm with $\epsilon=0.01$ on the Walker2d-Medium dataset, the threshold was too small to find any reachable states. In these cases, the reachable states were the explicit next state in the dataset, therefore Eq. equation 12 and equation 14 where only updated using $(s,s^{\prime})\in\mathcal{D}$. This is essentially the same method as TD3+BC, except using $Q(s,s^{\prime})$ instead of $Q(s,a)$. In these cases, StaCQ performs on par with TD3+BC, as expected, but much worse than StaCQ with a larger threshold.
Otherwise, for all metrics increasing the threshold leads to an increased score up to a certain point (e.g. $\epsilon=0.1$ for $L^{2}$-norm) where too many states are considered reachable, when they are not, at which point the average score decreases. On the Walker2d-medium dataset, StaCQ seems fairly robust to a change in reachability as even with the large thresholds the scores do not decrease as far as TD3+BC. On the Walker2d-MediumExpert dataset, when the threshold is too large, $\epsilon=0.5$ for the $L^{2}$ and $L^{\infty}$norms, the performance degrades massively. This is because too many states are considered reachable which are in fact not.
As a result of this study, the reachability metric chosen across all datasets is the $L^{\infty}$-norm with $\epsilon=0.1$, as these values produced the most reliable scores. It should be noted that other norms and thresholds could have been used to achieve similar or greater scores, however across all datasets these hyperparameters performed well.
Norm Threshold ($\epsilon$) Walker2d-Medium Walker2d-MediumExpert
$L^{2}$-norm $0.01$ $94.1\pm 3.1$ $116.0\pm 1.8$
$0.05$ $94.4\pm 1.4$ $104.5\pm 0.1$
$0.1$ $92.1\pm 1.0$ $94.0\pm 22.8$
$0.5$ $91.9\pm 2.5$ $0.3\pm 0.6$
$L^{1}$-norm $0.01$ $82.6\pm 1.3$ $115.5\pm 1.9$
$0.05$ $82.6\pm 1.3$ $114.6\pm 1.8$
$0.1$ $84.1\pm 17.0$ $113.7\pm 5.0$
$0.5$ $90.3\pm 2.5$ $107.8\pm 11.1$
$L^{\infty}$-norm $0.01$ $82.6\pm 1.3$ $115.7\pm 1.3$
$0.05$ $92.6\pm 3.6$ $114.2\pm 1.9$
$0.1$ $92.2\pm 3.0$ $116.2\pm 1.5$
$0.5$ $91.4\pm 5.8$ $6.7\pm 12.4$
Table 3: Comparison of different reachability metric norms and thresholds on the Walker2d-Medium and Walker2d-MediumExpert datasets. The normalised score is averaged for 10 evaluations providing a standard deviation over 5 seeds.
Appendix G Learning curves for StaCQ
In this section we provide the learning curves of StaCQ, across all datasets. During training, at 5000 gradient step intervals, the mean of 10 evaluations (for the locomotion tasks) or 100 evaluations (for the Antmaze tasks) of StaCQ had been recorded across 5 seeds. This is used to monitor the performance of StaCQ during training and the environment interactions are not used for the policy or Q-value updates.
(a) Hopper
(b) Walker2d
(c) Halfcheetah
(d) Antmaze-Umaze
(e) Antmaze-Medium
(f) Antmaze-Large
Figure 3: Learning curves for StaCQ across all datasets, grouped together by environment. The mean normalised score, over 10 evaluations (for locomotion tasks) or 100 evaluations (for Antmaze tasks), is recorded for every 5000 steps during training. The solid line is the mean score across 5 seeds and the shaded region is the standard deviation.
Appendix H Multi-step reachability: potential extensions and an illustrative example
The SCQL framework leverages one-step state reachability. A natural and promising direction for future work involves extending this concept to multi-step reachability. The core idea is to use a learned forward dynamics model, $f_{\omega_{1}}$, to simulate $k$-step trajectories $(s_{0},a_{0},\dots,a_{k-1},s_{k})$ starting from a state $s_{0}$. The crucial constraint remains that the trajectory must terminate at a state $s_{k}$ that belongs to the original dataset $\mathcal{D}$. This allows the agent to potentially bridge larger gaps in the dataset by traversing through intermediate states ($s_{1},\dots,s_{k-1}$) that might be out-of-distribution, while still anchoring the final value estimate to reliable in-distribution data ($s_{k}\in\mathcal{D}$).
The two-step case
Let us examine the specific case of two-step reachability. We define the set of 2-step reachable states ending in the dataset $\mathcal{D}$ from a state $s$ as:
$\mathcal{SR}^{(2)}_{\mathcal{M},\mathcal{D}}(s)=\{s_{2}\in\mathcal{D}\mid\exists a_{0},a_{1}\text{ s.t. }p(s_{1}|s,a_{0})=1\text{ and }p(s_{2}|s_{1},a_{1})=1\}.$ (34)
Here, the intermediate state $s_{1}$ does not need to be in $\mathcal{D}$. Practically, reachability would be estimated using the learned model $f_{\omega_{1}}$, denoted $\widehat{\mathcal{SR}}^{(2)}_{\mathcal{M},\mathcal{D}}(s)$.
To utilise this, a modified Q-learning update could estimate the value $Q(s,s^{\prime})$ based on the best 2-step reachable state from $s^{\prime}$. A conceptual sketch, analogous to Eq. equation 6, might look like:
$Q(s,s^{\prime})\leftarrow(1-\alpha)Q(s,s^{\prime})+\alpha\Bigl{[}r(s,s^{\prime})+\gamma\max_{\begin{subarray}{c}s^{\prime\prime\prime}\in\mathcal{D}\\ \cap\\ s^{\prime\prime\prime}\in\widehat{\mathcal{SR}}^{(2)}_{\mathcal{M},\mathcal{D}}(s^{\prime})\end{subarray}}V^{(2)}(s^{\prime},s^{\prime\prime\prime})\Bigr{]}.$ (35)
In this sketch, $V^{(2)}(s^{\prime},s^{\prime\prime\prime})$ represents the estimated value of the optimal 2-step path initiating from $s^{\prime}$ and terminating at $s^{\prime\prime\prime}\in\mathcal{D}$. Accurately estimating $V^{(2)}$ using the learned model $f_{\omega_{1}}$ (involving rewards for the first step $s^{\prime}\to s^{\prime\prime}$ and the value/reward for the second step $s^{\prime\prime}\to s^{\prime\prime\prime}$) is a key challenge, requiring careful handling of model predictions and potential accumulated errors.
Extracting an optimal policy $\pi^{*}(s)$ based on $k$-step reachability is also more complex than the 1-step case. Conceptually, the goal is to select the first action $a_{0}$ that initiates the sequence leading to the highest-value $k$-step reachable state $s_{k}^{*}\in\mathcal{D}$:
  1. 1.
    Find the best $k$-step endpoint: $s_{k}^{*}=\operatorname*{arg\,max}_{s_{k}\in\widehat{\mathcal{SR}}^{(k)}_{\mathcal{M},\mathcal{D}}(s)}V^{(k)}(s,s_{k})$.
  2. 2.
    Determine the first action $a_{0}^{*}$ on the optimal path from $s$ to $s_{k}^{*}$.
Step 2 typically involves some form of planning (e.g., sampling action sequences with the model) or learning a policy specifically optimised for this multi-step objective.
Illustration: maze two-step reachability
The potential benefit of multi-step reachability is illustrated in the simple maze environment shown in Figure 4. The dataset provided (Figure 4a) is sparse, preventing the standard one-step SCQL from finding a path to the goal from all dataset states (Figure 4b). However, by considering two-step reachability ($k=2$), the agent can plan through an intermediate state that may not be in the dataset (indicated by dotted arrows). Assuming perfect knowledge of the dynamics for this illustrative example, a hypothetical two-step state-constrained policy can successfully connect all starting points to the goal region (Figure 4c), demonstrating improved trajectory stitching capabilities.
(a) Maze and Dataset
(b) SCQL: 1-step reachable policy
(c) SCQL: 2-step reachable policy
Figure 4: Comparison of one-step and two-step state reachability SCQL methods on a simple maze environment. (a) The maze is a 10 by 10 grid where the coordinate values $(x,y)$ represent the state. The high reward region is represented by a star and maze walls are represented by red grid squares. The dataset is composed of 4 trajectories represented by the dotted lines where the white circle is the starting state and the black circle is the final state. (b) The final policy when applying one-step reachable SCQL to the dataset. (c) The final policy when applying two-step reachable SCQL to the dataset, where dotted arrows represent transitions through a state that is not in the dataset.
Potential challenges
While promising, extending state-constrained RL to multiple steps also introduces significant practical challenges, primarily stemming from reliance on the learned dynamics model $f_{\omega_{1}}$. Key challenges and potential research directions include:
  • Model error accumulation: Errors in $f_{\omega_{1}}$ compound over longer rollouts ($k>1$), leading to inaccurate state predictions and potentially unreliable value estimates for the $k$-step paths.
  • Handling uncertainty: Effectively leveraging multi-step rollouts requires managing this model uncertainty. Potential approaches include incorporating pessimism into value estimates based on rollout uncertainty (Yu et al., 2020; Rigter et al., 2022), using model ensembles to quantify confidence, or dynamically adapting the rollout horizon $k$ based on model reliability.
  • Computational cost: Finding and evaluating all $k$-step reachable paths terminating in $\mathcal{D}$ can be computationally demanding. Efficient search or sampling techniques would be crucial. Graph-based search algorithms operating on the state space defined by the learned model could offer an alternative to simple rollouts for finding paths back to $\mathcal{D}$.
  • Algorithmic integration: Developing robust Q-learning or policy optimization methods that correctly incorporate these multi-step, state-constrained value estimates is non-trivial, especially when considering the complexities of planning or differentiating through model rollouts.
Exploring these multi-step extensions involves balancing the benefit of connecting more distant states against the cost and unreliability of longer model-based rollouts. This intersection of model-based planning and offline constraints presents a rich area for future investigation within the state-constrained RL paradigm.